one number is eight more than another. their product is -16. what are the numbers?

Answer by Boreal(15235)

numbers are x, and x+8
product is -16
x(x+8)=x^2+8x
it equals -16
x^2+8x=-16
x^2+8x+16=0
(x+4)^2=0
x=-4
x+8=4
numbers are -4 and 4.